OpenStudy (anonymous):

The container that holds the water for the football team is 1/2 full. After pouring in 3 gallons of water, it is 2/3 full. How many gallons can the container hold?

OpenStudy (anonymous):

Let x be the gallon capacity of the container. Solve the following for x: \[\frac{1}{2}x+3=\frac{2}{3}x\]\[x=18 \text{ gallons}\]

OpenStudy (anonymous):

Thank you.